OpenStudy (anonymous):

find the domain of the function f(X) = ln(2-x)

OpenStudy (anonymous):

domain of this function would be x<=2 since log is not defined for negative values

OpenStudy (anonymous):

so it is (- infinity, 2]

OpenStudy (anonymous):

got it jsy? or u need more clarification? :):)

OpenStudy (anonymous):

sry *jay

OpenStudy (anonymous):

Are you there???? @jay

OpenStudy (amistre64):

ln is not define for zero either....when does ln(x) = 0?

OpenStudy (anonymous):

log 0 = 1 am i right?

OpenStudy (anonymous):

if its so, then the answer is (-infinity, 2)

OpenStudy (amistre64):

no...sorry :) thats better.. think of it this way. the log points to an exponent: log2(8) = 3 the exponent is "3" the base is to "2" 2^3 = 8 now tell me.... 2^? = 0 ??

OpenStudy (amistre64):

log2(0) = ?

OpenStudy (anonymous):

then my second option is correct? isn't it? and not the first one...

OpenStudy (amistre64):

(-inf,2) is correct :)

OpenStudy (anonymous):

hmm.... yay amistre:)

OpenStudy (anonymous):

i think jay is not here

OpenStudy (amistre64):

jay was just the catalyst.... math goes on without him lol

OpenStudy (anonymous):

:)but i think this shud help him

OpenStudy (anonymous):

i definitely see that as an option thank u so much

OpenStudy (anonymous):

you 're welcome:) any more questions??

OpenStudy (anonymous):

\[2-x \ge0\]\[x \le 2\] hence the domain is \[[-\infty, 2)\]